# Loop-Residual Neural Networks for Iterative Refinement

### v2

@@ Abstract
The success of large-scale language models like GPT can be attributed to their ability to efficiently predict the next token in a sequence. However, these models rely on constant computational effort regardless of the complexity of the token they are predicting, lacking the capacity for iterative refinement. In this paper, we introduce a novel Loop-Residual Neural Network, which achieves better performance by utilizing longer computational time without increasing the model size. Our approach revisits the input multiple times, refining the prediction by iteratively looping over a subset of the model with residual connections. We demonstrate the effectiveness of this method through experiments comparing versions of GPT-2 with our Loop-Residual models, showing improved performance in language modeling tasks while maintaining similar parameter counts. Importantly, these improvements are achieved without the need for extra training data.
@@ 1 Introduction
In this paper, we introduce a novel Loop-Residual Neural Network, which revisits the input multiple times, refining the prediction by iteratively looping over a subset of the model with residual connections. Our main contribution is improving transformer performance with longer inference times, using a novel loop architecture with residual prediction. This approach works for large neural networks without needing extra training data, effectively extending the model’s approximation capacity.
@@
  • Novel Architecture: We introduce a loop-residual mechanism that enhances model performance without increasing parameter count, making it accessible for researchers with limited computational resources.
@@
We demonstrate the effectiveness of this method through experiments comparing versions of GPT-2 with our Loop-Residual models, showing that our GPT-2-81M model achieves a validation loss of 3.11 on the OpenWebText dataset (Gokaslan and Cohen,, 2019), comparable to the GPT-2-124M model’s loss of 3.12. These results emphasize the benefit of utilizing longer inference times to improve performance without increasing model size or requiring additional data.
@@ 2 Background and Related Work
Residual connections, introduced by He et al., (2016), have been instrumental in enabling the training of very deep networks by alleviating the vanishing gradient problem. Our Loop-Residual model leverages this concept by focusing on refining the hidden state through iterative loops, allowing the model to stabilize training and accelerate convergence without increasing its size significantly.
2.1 Iterative Refinement and Adaptive Computation
Several prior works have explored iterative refinement and adaptive computation in neural networks.
@@ 2.1.3 Depth-Adaptive Transformers
Elbayad et al., (2019) presented Depth-Adaptive Transformers that adjust the depth of the network based on the input. Their method enables dynamic inference by selecting the number of layers to apply per input sequence. However, their approach focuses on varying depth during inference rather than iterative refinement during training, and their experiments were conducted on machine translation tasks with relatively small models. Importantly, their model lacks the predictive residual design present in our Loop-Residual architecture.
@@ 3 Methodology
We propose a novel architecture called the Loop-Residual Neural Network, which refines the hidden state through iterative loops over a subset of transformer blocks with residual connections and gating mechanisms. This iterative process allows the model to revisit and enhance its internal representations multiple times, effectively increasing its approximation capacity without adding more parameters.
@@ 3.1 General Loop Structure with Residual Design
3.2 Illustration of the Loop-Residual Mechanism
@@ 4 Experiments
We conducted experiments to evaluate the effectiveness of our Loop-Residual Neural Network in improving transformer performance using longer inference times without increasing model size. Our experiments focus on comparing our models with standard GPT-2 models of similar or larger sizes on the OpenWebText dataset.
@@ 4.1 Experimental Setup
In our first experiment, we compared a Loop-Residual GPT-2 model with 81 million parameters (GPT2-81M) to the standard GPT-2 model with 124 million parameters (GPT2-124M). The GPT2-124M model consists of 12 transformer layers, serving as our baseline. Our Loop-Residual GPT2-81M model employs 6 loops over 6 transformer layers, effectively simulating deeper computation without increasing the number of parameters. The objective of this experiment is to demonstrate that our model achieves comparable or better performance using fewer parameters by leveraging longer inference times through iterative refinement.
In our second experiment, we focused on smaller-scale models to illustrate the effectiveness of our method at different scales. We compared a Loop-Residual GPT-2 model with 45 million parameters (GPT2-45M) to a Lite GPT-2 model of the same size (GPT2-45M-Lite). The GPT2-45M-Lite model consists of a single transformer block layer, performing one-pass prediction, and serves as the baseline. Our Loop-Residual GPT2-45M model loops twice over a single transformer block to predict the residual. The objective here is to illustrate that even at smaller scales, our method can improve performance over the baseline without increasing model size by leveraging iterative refinement.
We selected these experimental settings to evaluate the efficacy of our loop-residual mechanism across different model sizes and to show that iterative refinement can compensate for a reduced number of parameters or layers. By keeping the total number of parameters constant while adjusting the depth via looping, we aim to isolate the effect of iterative refinement on model performance.
In all experiments, we used the same training data and hyperparameters where applicable to ensure a fair comparison. The models were trained on the OpenWebText dataset, which provides a diverse range of textual data suitable for language modeling tasks. We measured the training epoch times for our models and the standard GPT-2 models to assess the computational overhead introduced by our method. The training epoch times were 150ms for GPT2-45M-Lite, 177ms for our Loop-Residual GPT2-45M, and 1,377ms for GPT2-81M.
@@ 5 Results
Our experiments demonstrate that the Loop-Residual Neural Network significantly improves performance over standard transformer models by leveraging longer inference times without increasing model size.
5.1 First Experiment: GPT2-81M vs. GPT2-124M
Our Loop-Residual GPT2-81M model achieved a validation loss of 3.11 on the OpenWebText dataset, comparable to the GPT2-124M model’s loss of 3.12. This result is notable because our model uses approximately 35% fewer parameters and half the number of unique layers compared to the GPT2-124M model. The improvement indicates that iterative refinement through our loop-residual mechanism effectively enhances the model’s approximation capacity, allowing it to match or exceed the performance of larger models.
@@
Model Parameters Layers Loops Validation Loss
@@
Loop-Residual GPT2-81M 81M 6 6 3.11
2020-20- 20002020202040404040606060608080808010010010010012012012012014014014014016016016016018018018018020020020020022022022022033334444555566667777888899991010101011111111Training Steps (in thousands)LossLoop-Residual GPT2-81M Train LossLoop-Residual GPT2-81M Val LossGPT2-124M Val Loss (12 Layers)
@@
5.2 Second Experiment: Loop-Residual GPT2-45M vs. Lite GPT2-45M
In the comparison between the Loop-Residual GPT2-45M model and the Lite GPT2-45M model, our model achieved a validation loss of 3.67 compared to 3.98, and a training loss of 3.65 compared to 3.96. By looping twice over a single transformer block, our model effectively simulates a deeper network, resulting in a substantial performance gain over the one-pass baseline. This demonstrates that even at smaller scales, our method can improve performance without increasing model size by leveraging iterative refinement.
@@
Lite GPT2-45M 45M 1 1 3.96 3.98
Loop-Residual GPT2-45M 45M 1 2 3.65 3.67
2020-20- 20002020202040404040606060608080808010010010010012012012012014014014014016016016016018018018018020020020020022022022022033334444555566667777888899991010101011111111Training StepsLossLoop-Residual GPT2-45M (1 Layer, 2 Loops) Train LossLoop-Residual GPT2-45M Val LossLite GPT2-45M (1 Layer) Train LossLite GPT2-45M Val Loss
@@ 5.3 Discussion
The results indicate that our Loop-Residual Neural Network architecture effectively leverages iterative refinement to enhance model performance without increasing the parameter count. In the first experiment, despite having approximately 35% fewer parameters and half the number of unique layers compared to the GPT2-124M baseline, our Loop-Residual GPT2-81M model achieves a slightly better validation loss. This suggests that the iterative loops over the transformer blocks allow the model to capture complex patterns and dependencies that would otherwise require additional layers or parameters.
In the second experiment, the Loop-Residual GPT2-45M model significantly outperforms the GPT2-45M-Lite baseline. By looping twice over a single transformer block, our model effectively simulates a deeper network, resulting in a reduction of the validation loss from 3.98 to 3.67. This demonstrates that even at smaller scales, iterative refinement can compensate for a limited number of layers, enhancing the model’s expressiveness.
The increased training epoch times for our Loop-Residual models are relatively modest compared to the performance gains. For instance, the training time for the Loop-Residual GPT2-45M model is only 18% longer than the GPT2-45M-Lite baseline, while achieving a substantial improvement in validation loss. This indicates that the computational overhead introduced by the loops is manageable, making our approach practical for applications where longer inference times are acceptable in exchange for improved performance.
Our experiments highlight the advantage of our setup in terms of resource efficiency and scalability. By avoiding the need for additional parameters or larger models, our Loop-Residual Neural Network is more accessible for researchers and practitioners with limited computational resources. Furthermore, since our approach does not rely on additional training data or specialized datasets, it offers a scalable solution for enhancing model performance in various language modeling tasks.
@@ 6 Conclusion
We have presented the Loop-Residual Neural Network, a novel approach that enables smaller neural network models to achieve better results on lower-end devices by leveraging longer inference times through iterative refinement. By iteratively refining the hidden state through residual connections over multiple loops of a model subset, our method captures complex patterns and dependencies more effectively than conventional one-pass models. Our experiments demonstrate that the Loop-Residual models can achieve improved performance over baseline models of the same size and comparable performance to larger models with fewer parameters. Significantly, these performance improvements are achieved without the need for additional training data, highlighting the efficiency and accessibility of our approach. This work opens up new possibilities for neural network architectures, particularly for tasks that benefit from deeper computational reasoning on resource-constrained devices.

### v3

# Loop Neural Networks for Parameter Sharing

@@ Abstract
The success of large-scale language models like GPT can be attributed to their ability to efficiently predict the next token in a sequence. However, these models rely on constant computational effort regardless of the complexity of the token they are predicting, lacking the capacity for iterative refinement. In this paper, we introduce a novel loop neural network, which achieves better performance by utilizing longer computational time without increasing the model size. Our approach revisits the input multiple times, refining the prediction by iteratively looping over a subset of the model with residual connections. We demonstrate the effectiveness of this method through experiments comparing versions of GPT-2 with our loop neural networks, showing improved performance in language modeling tasks while maintaining similar parameter counts. Importantly, these improvements are achieved without the need for extra training data.
@@ 1 Introduction
In this paper, we introduce a novel loop neural network, which revisits the input multiple times, refining the prediction by iteratively looping over a subset of the model with residual connections. Our main contribution is improving transformer performance with longer inference times, using a novel loop architecture with residual prediction. This approach works for large neural networks without needing extra training data, effectively extending the model’s approximation capacity.
@@
  • Novel Architecture: We introduce a loop neural network mechanism that enhances model performance without increasing parameter count, making it accessible for researchers with limited computational resources.
@@
We demonstrate the effectiveness of this method through experiments comparing versions of GPT-2 with our loop neural networks, showing that our loop neural network with 6 transformer layers achieves a validation loss of 3.11 on the OpenWebText dataset (Gokaslan and Cohen,, 2019), comparable to the GPT2 model’s loss of 3.12 with 12 transformer layers . Similarly, our loop neural network with 4 layers and 12 loops achieves a validation loss of 3.15, only less than 1% increase compared to the GPT2-124M model. These results emphasize the benefit of utilizing longer inference times to improve performance without increasing model size or requiring additional data.
@@ 2 Background and Related Work
Residual connections, introduced by He et al., (2016), have been instrumental in enabling the training of very deep networks by alleviating the vanishing gradient problem. Our loop neural network leverages this concept by focusing on refining the hidden state through iterative loops, allowing the model to stabilize training and accelerate convergence without increasing its size significantly.
2.1 Parameter Sharing and Adaptive Computation
Several prior works have explored parameter sharing and adaptive computation in neural networks.
@@ 2.1.3 Depth-Adaptive Transformers
Elbayad et al., (2019) presented Depth-Adaptive Transformers that adjust the depth of the network based on the input. Their method enables dynamic inference by selecting the number of layers to apply per input sequence. However, their approach focuses on varying depth during inference rather than iterative refinement during training, and their experiments were conducted on machine translation tasks with relatively small models. Importantly, their model lacks the predictive residual design present in our loop architecture.
2.1.4 Parameter Sharing Across Layers
Parameter sharing across layers has been explored as a way to reduce the number of parameters in transformer models while maintaining performance. Takase and Kiyono, (2023) investigated the effects of parameter sharing in transformers, showing that sharing parameters across layers can be effective under certain conditions. In our work, we employ a further idea by looping over layers with shared parameters with residual and gate design, effectively increasing computational depth without adding new parameters.
@@ 3 Methodology
We propose a novel architecture called the loop neural network, which refines the hidden state through iterative loops over a subset of transformer blocks with residual connections and gating mechanisms. This iterative process allows the model to revisit and enhance its internal representations multiple times, effectively increasing its approximation capacity without adding more parameters.
@@ 3.1 General Loop Structure with Residual Design
3.2 Looping Strategies
Our approach applies the loop in two ways: looping over all layers multiple times, and looping within each layer. This strategy effectively increases the computational depth of the model without increasing the parameter count, similar to techniques explored in parameter sharing across transformer layers.
3.3 Illustration of the Loop Neural Network
@@ 4 Experiments
We conducted experiments to evaluate the effectiveness of our loop neural networks in improving transformer performance using longer inference times without increasing model size. Our experiments focus on comparing our models with standard GPT-2 models of similar or larger sizes on the OpenWebText dataset.
@@ 4.1 Experimental Setup
In our first experiment, we compared loop neural networks with different configurations to the standard GPT-2 models. Specifically, we evaluated:
  • GPT2-124M: The standard GPT-2 model with 124 million parameters and 12 layers.
  • GPT2-81M-LOOP: Our loop neural network with 81 million parameters, 6 layers, and 6 loops.
  • GPT2-67M-LOOP: Our loop neural network with 67 million parameters, 4 layers, and 12 loops.
All other structural and design elements of the model, such as the transformer block configurations and parameter initialization, remain identical to the baseline GPT-2 models, with the only modifications being the loop residual and gating mechanisms. The GPT2-124M model serves as our baseline. Our GPT2-81M-LOOP model employs 6 loops over 6 transformer layers, effectively simulating deeper computation without increasing the number of parameters. The GPT2-67M-LOOP model uses 6 loops over 4 transformer layers and 2 loops over each layer. This model demonstrates that even with fewer layers and parameters, we can achieve comparable performance by increasing the number of loops.
In our second experiment, we focused on smaller-scale models to illustrate the effectiveness of our method at different scales. We compared a loop neural network with 45 million parameters (GPT2-45M-LOOP) to a GPT2-45M model of the same size. The GPT2-45M model consists of a single transformer block layer, performing one-pass prediction, and serves as the baseline. Our GPT2-45M-LOOP model loops twice over a single transformer block to predict the residual. The objective here is to illustrate that even at smaller scales, our method can improve performance over the baseline without increasing model size by leveraging iterative refinement.
We selected these experimental settings to evaluate the efficacy of our loop neural network across different model sizes and to show that iterative refinement can compensate for a reduced number of parameters or layers. By keeping the total number of parameters constant while adjusting the depth via looping, we aim to isolate the effect of iterative refinement on model performance.
In all experiments, we used the same training data and hyperparameters where applicable to ensure a fair comparison. The models were trained on the OpenWebText dataset, which provides a diverse range of textual data suitable for language modeling tasks. We measured the training epoch times for our models and the standard GPT-2 models to assess the computational overhead introduced by our method. The training epoch times were 150ms for GPT2-45M, 177ms for our GPT2-45M-LOOP, and 1,377ms for GPT2-81M-LOOP.
@@ 5 Results
Our experiments demonstrate that the loop neural networks significantly improve performance over standard transformer models by leveraging longer inference times without increasing model size.
5.1 First Experiment: Loop Neural Networks vs. GPT2-124M
Our GPT2-81M-LOOP model achieved a validation loss of 3.11 on the OpenWebText dataset, comparable to the GPT2-124M model’s loss of 3.12. Similarly, the GPT2-67M-LOOP model, with only 67 million parameters and 4 layers, achieved a validation loss of 3.15, only less than 1% increase compared to the GPT2-124M model. These results are notable because our models use significantly fewer parameters and layers compared to the GPT2-124M model. The improvement indicates that iterative refinement through our looping mechanism effectively enhances the model’s approximation capacity, allowing it to match or exceed the performance of larger models.
@@
Model Parameters Layers Loops Train Loss Validation Loss
@@
GPT2-81M-LOOP 81M 6 6 3.13 3.11
GPT2-67M-LOOP 67M 4 12 3.15 3.15
2020-20- 20002020202040404040606060608080808010010010010012012012012014014014014016016016016018018018018020020020020022022022022033334444555566667777888899991010101011111111Training Steps (in thousands)LossGPT2-81M-LOOP Train LossGPT2-81M-LOOP Val LossGPT2-124M Val Loss (12 Layers)
@@
5.2 Second Experiment: GPT2-45M-LOOP vs. GPT2-45M
In the comparison between the GPT2-45M-LOOP model and the GPT2-45M model, our model achieved a validation loss of 3.67 compared to 3.98, and a training loss of 3.65 compared to 3.96. By looping twice over a single transformer block, our model effectively simulates a deeper network, resulting in a substantial performance gain over the one-pass baseline. This demonstrates that even at smaller scales, our method can improve performance without increasing model size by leveraging iterative refinement.
@@
GPT2-45M 45M 1 1 3.96 3.98
GPT2-45M-LOOP 45M 1 2 3.65 3.67
2020-20- 20002020202040404040606060608080808010010010010012012012012014014014014016016016016018018018018020020020020022022022022033334444555566667777888899991010101011111111Training StepsLossGPT2-45M-LOOP (2 Loops) Train LossGPT2-45M-LOOP Val LossGPT2-45M Train LossGPT2-45M Val Loss
@@ 5.3 Discussion
The results indicate that our loop architecture effectively leverages iterative refinement to enhance model performance without increasing the parameter count. In the first experiment, despite having significantly fewer parameters and layers compared to the GPT2-124M baseline, our GPT2-81M-LOOP and GPT2-67M-LOOP models achieve comparable validation losses. This suggests that the iterative loops over the transformer blocks allow the models to capture complex patterns and dependencies that would otherwise require additional layers or parameters.
In the second experiment, the GPT2-45M-LOOP model significantly outperforms the GPT2-45M baseline. By looping twice over a single transformer block, our model effectively simulates a deeper network, resulting in a reduction of the validation loss from 3.98 to 3.67. This demonstrates that even at smaller scales, iterative refinement can compensate for a limited number of layers, enhancing the model’s expressiveness.
The increased training epoch times for our loop neural networks are relatively modest compared to the performance gains. For instance, the training time for the GPT2-45M-LOOP model is only 18% longer than the GPT2-45M baseline, while achieving a substantial improvement in validation loss. This indicates that the computational overhead introduced by the loops is manageable, making our approach practical for applications where longer inference times are acceptable in exchange for improved performance.
Our experiments highlight the advantage of our setup in terms of resource efficiency and scalability. By avoiding the need for additional parameters or larger models, our loop neural network is more accessible for researchers and practitioners with limited computational resources. Furthermore, since our approach does not rely on additional training data or specialized datasets, it offers a scalable solution for enhancing model performance in various language modeling tasks.
@@ 6 Conclusion
We have presented the loop neural network, a novel approach that enables smaller neural network models to achieve better results on lower-end devices by leveraging longer inference times through iterative refinement. By iteratively refining the hidden state through residual connections over multiple loops of a model subset, our method captures complex patterns and dependencies more effectively than conventional one-pass models. Our experiments demonstrate that the loop neural networks can achieve improved performance over baseline models of the same size and comparable performance to larger models with fewer parameters. Significantly, these performance improvements are achieved without the need for additional training data, highlighting the efficiency and accessibility of our approach. This work opens up new possibilities for neural network architectures, particularly for tasks that benefit from deeper computational reasoning on resource-constrained devices.